OpenStudy (anonymous):

If x-y=10 And x^2-y^2=40 then y= a)14 b)7 c)4 d)-3 e)-6

OpenStudy (anonymous):

Plug them in: x-14=10 x=24 24^2-14^2=40? Nope. x-7=10 x=17 17^2-7^2=40? Nope. x-4=10 x=14 14^2-4^2=40? Nope. x-(-3)=10 x+3=10 x=7 7^2-(-3)^2=40? Yup. x-(-6)=10 x+6=10 x=4 (-6)^2-4^2=40? Nope.